a) 35=5x?
b)50=2x5x ?
c)30=2x?x?
d)54=2x3
1
Expert's answer
2019-07-09T11:53:16-0400

a) 35 = 5x,

x = 7.

b) 50 = 2x*5x,

10x^2 = 50,

x^2 = 5,

x1 = 2.24, x2 = -2.24.

c) 30 = 2x*x,

x^2 = 15,

x1 = 3.87, x2 = -3.87.

d) 54 = 2x^3,

x^3 = 27,

x = 3.